OpenStudy (anonymous):

lim as x --> 0 ((sin3x)(sin2x))/(xsin5x)

OpenStudy (anonymous):

still no l'hopital, lots of algebra. answer is \[\frac{6}{5}\] myininaya will write it out

OpenStudy (anonymous):

i am anxiously awaiting this algebra. also a reasonable explanation of how you decide what to do

OpenStudy (anonymous):

likewise! thanks for the help.

OpenStudy (anonymous):

(only saying that because i have no idea how to do it)

myininaya (myininaya):

first of all we know \[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1 \] by squeeze thm now let's prove something else \[\lim_{x \rightarrow 0}\frac{\sin(Ax)}{\sin(Bx)}=\frac{A}{B}\] \[\lim_{x \rightarrow 0}\frac{\sin(Ax)}{\sin(Bx)} \cdot \frac{ABx}{ABx}=\lim_{x \rightarrow 0}\frac{A}{B} \cdot \frac{\sin(Ax)}{Ax} \cdot \frac{Bx}{\sin(Bx)}=\frac{A}{B} \cdot 1 \cdot 1=\frac{A}{B}\] so we have \[\lim_{x \rightarrow 0}\frac{\sin(3x)}{x} \cdot \frac{\sin(2x)}{\sin(5x)} \cdot \frac{3}{3}\] \[=3\lim_{x \rightarrow 0}\frac{\sin(3x)}{3x} \cdot \frac{\sin(2x)}{\sin(5x)}=3 \cdot 1 \cdot \frac{2}{5}\]

OpenStudy (anonymous):

very nice explanation!

myininaya (myininaya):

:)

OpenStudy (anonymous):

awesome. you blew my mind!

OpenStudy (anonymous):

myininaya is the limit!